OpenStudy (anonymous):

algebra help please how do you solve 8y^2+12+4=y+10??

OpenStudy (anonymous):

12y

sam (.sam.):

You sure you wrote it correctly?

sam (.sam.):

you'll get imaginary number

OpenStudy (anonymous):

8y^2+12y+4+y+10 it has to do with factoring polynomials

sam (.sam.):

Fator? 8y^(2)+12y+4+y+10 8y^(2)+13y+14 Can't factor further

OpenStudy (anonymous):

nooo my apologies 4=Y+10

sam (.sam.):

but even though its 8y^2+12+4=y+10, you'll get imaginary number

OpenStudy (anonymous):

imaginary numbers with (i) you will get it